OpenStudy (anonymous):

What do you need to add to 3m^3 +7m +12 to get the sum of 2m^2 + 3m- 7 ?

OpenStudy (saifoo.khan):

Yes, exactly.

OpenStudy (anonymous):

2m^2 + 3m- 7 - (3m^3 +7m +12)

OpenStudy (anonymous):

Could you please show me the process

OpenStudy (anikate):

I can do this just give me some time

OpenStudy (anonymous):

distribute the minus sign, then collect like terms. what do you get?

OpenStudy (anikate):

It's long an d confusing

OpenStudy (anikate):

Yes first distribute minus then

OpenStudy (anonymous):

collect like terms

OpenStudy (anikate):

then add like terms

OpenStudy (saifoo.khan):

3m^3 +7m +12 to get the sum of 2m^2 + 3m- 7 Arrange, 3m^3 + 2m^2 + 3m + 7m + 12 - 7 Now add terms with same variables and their expoents. 3m^3 + 2m^2 + 10m + 5

OpenStudy (anikate):

Cool Saif you are a legend

OpenStudy (saifoo.khan):

Thanks Anikate! =)

OpenStudy (anikate):

yw

OpenStudy (anikate):

np

OpenStudy (anonymous):

Yes but when we add 3m^3 and 3m^3, it becomes 6m^3, but there is not m^3 in the answer

OpenStudy (saifoo.khan):

u didnt wrote any 3m^3 in the second thing.

OpenStudy (anonymous):

No, the question goes like this: 3m^3 +7m +12 + (BLANK) = 2m^2 + 3m- 7 ?

OpenStudy (anonymous):

-3m^3 + 2m^2 - 4m -19

OpenStudy (saifoo.khan):

What money bird said..

OpenStudy (saifoo.khan):

Actually u have to solve it this way, like u have: 2 + x = 5 And u have to find that "x" So for that u have to solve for x. x = 5 - 2 x =3

OpenStudy (anonymous):

And moneybird how did you get that?

OpenStudy (anonymous):

2m^2 + 3m- 7 - (3m^3 +7m +12)

OpenStudy (saifoo.khan):

Now, Apply the same example here, 3m^3 +7m +12 + (BLANK) = 2m^2 + 3m- 7 U have to solve for the "blank" "Blank" = (2m^2 + 3m - 7) - (3m^3 + 7m + 12)

OpenStudy (mimi_x3):

unfathomable^

OpenStudy (anonymous):

so the supposed answer minus the first question

OpenStudy (saifoo.khan):

@Mini stop spamming and get a life.

OpenStudy (mimi_x3):

im boredddddddddddd

OpenStudy (saifoo.khan):

Yes! this is b/c as u are going to shift the expression on the other side, it will change it's signs.